# Exact Bayesian Inference for Markov Switching Diffusions

### v1

@@ 3.4 Generator Update
Since the full conditional density for $\Lambda$ does not involve the augmented transition densities, it is the easiest update to carry out:
$\pi(\lambda|y)\propto\pi(y|\lambda)\pi(\lambda).$ (37)
For jump times $r$, the density $\pi(y|\lambda)$ with respect to $\mathbb{L}$ is given by
@@
$\pi(y|\lambda)\propto\prod_{i}\left\lparen e^{-\lambda_{i}\chi_{i}}\prod_{j\neq i}\lambda_{ij}^{n_{ij}}\right\rparen.$ ) . (39)
@@ 3.5 Practical Considerations
In diffusion settings, the main practical challenge that Bernoulli factory MCMC face is proposal sensitivity - the notion that the 2-coin algorithm runtime can explode when proposing a move to some regions of the state space, especially those associated with large diffusion drift. This is due to the fact that in such regions, the discrepancy between the diffusion measure and the Brownian bridge proposals is large, and the success probability of Poisson coins is low. Under a correctly specified model, the diffusion path avoids regions of $\mathcal{V}$ with high drift, and conversely, parameters associated with large drift usually have low posterior probability. Accordingly, the $\theta$-proposals in particular need to be carefully chosen. We adopt two mitigations to this issue, the first of which is to apply the Portkey Barker algorithm of (vats2020efficient), which truncates the infinite 2-coin loop at the cost of lesser MCMC efficiency. In practice, the main advantage of this modification is that it discards proposals with potentially very low acceptance probability more cheaply. Moreover, the Portkey approach can be extended to the divide-and-conquer Bernoulli factory. The other mitigation consists of initializing the chain and pre-adapting the proposals by running a chain on an Euler-approximated posterior first. We specify the approximation in Supplement C. This allows us to then initialize the exact chain at a better location in the state space, with reasonable tuning parameters. We adopt both mitigation strategies in Sections 5 and 6.
More specific challenges apply to Markov switching diffusions, which can exhibit large drift even under the correct specification. This occurs e.g. when switching between two regimes with different stationary means, upon which the process may experience strong drift towards the new equilibrium. This is partially mitigated by refining the bounds on the diffusion path, as recommended by gonccalves2023exact, which increases the success probability of Poisson coins. The shortcoming of this intervention is the difficulty of tuning it adaptively during the MCMC run, requiring a full reset of the chain to modify the extent of the refinement. We also note that fairly strong posterior dependence between $\Theta$ and $Y$ can arise, which negatively affects mixing of the Gibbs chain. This tendency, while not catastrophic, is observed in the experiment in Section 5. Indeed, the corresponding model is invariant to label permutations, and the Markov chain will typically only visit one of the posterior’s equivalent modes. Moreover, it is common for the chain to drop one of the states during the transient phase, upon which the update for the parameters corresponding to that state may walk randomly. Therefore, during adaptation, we set the parameters of any inactive state equal to the parameters of an active state, which results in quick re-introduction of the inactive state.
@@ 4 Exact MAP and Maximum Likelihood Estimation
$(\theta^{\ddagger},\lambda^{\ddagger})=\operatorname*{\mathrm{argmax}}_{\theta,\lambda}\ \pi(\theta,\lambda,v_{s\mathbin{\backslash}\{0\}}|v_{0}).$ (43)
The MAP estimator also corresponds to the maximum likelihood estimator when setting $\pi(\theta,\lambda)\propto 1$. In this section, we adapt an approach originally proposed for Ito diffusions in beskos2006exact. It consists of constructing a Monte Carlo EM algorithm, which alternates between an approximate E-step, where we construct a Monte Carlo estimator $\bar{Q}$ of the function
$Q(\theta^{\dagger},\lambda^{\dagger},\theta,\lambda)=\operatorname{\mathrm{E}}_{V_{R},Z,Y}\left[\log\pi(V_{R},Z,Y,\theta^{\dagger},\lambda^{\dagger},v_{s\mathbin{\backslash}\{0\}}|v_{0})|v_{s},\theta,\lambda\right],$ (44)
@@
One benefit of the MCEM approach is that there is a substantial methodological overlap with posterior sampling, and we can make use of the hidden data update of Section 3.2 in devising the E-step. The MCEM algorithm also exploits the same conditional independence structure. Moreover, because it merely requires an unbiased estimate of the log-likelihood, the MCEM algorithm avoids some of the scaling issues with the MCMC parameter update, discussed in Section 3.3. Conversely, even an “exact” MCEM algorithm is not guaranteed to converge to the global maximum, though convergence to a local maximum can be proven in some circumstances (fort2003convergence).
We discuss the construction of $\bar{Q}$ in Section 4.1 before proceeding with the M-step in Section 4.2, and close out with a discussion of various implementation aspects in Section 4.3.
@@ 4.1 E-Step
In standard EM algorithms, the E-step consists of finding a lower bound on the objective $\pi(\theta,\lambda,v_{s\mathbin{\backslash}\{0\}}|v_{0})$. It is obtained by averaging the joint density over the posterior of the latent variables, i.e.
$\begin{split}Q(\theta^{\dagger},\lambda^{\dagger},\theta,\lambda)&=\operatorname{\mathrm{E}}_{V_{R},Z,Y}\left[\log\pi(V_{R},Z,Y,\theta^{\dagger},\lambda^{\dagger},v_{s\mathbin{\backslash}\{0\}}|v_{0})|v_{s},\theta,\lambda\right]\\ &=\operatorname{\mathrm{E}}_{V_{R},Z,Y}\left[\log\pi(V_{R},Z,v_{s\mathbin{\backslash}\{0\}}|Y,\theta^{\dagger},v_{0})+\log\pi(Y|\lambda^{\dagger})|v_{s},\theta,\lambda\right]\\ &\qquad+\log\pi(\theta^{\dagger})+\log\pi(\lambda^{\dagger}).\end{split}$ (46)
@@
$Q_{\Lambda}(\lambda^{\dagger},\lambda)=\operatorname{\mathrm{E}}_{V_{R},Z,Y}\left[\log\pi(Y|\lambda^{\dagger})|v_{s},\theta,\lambda\right]+\log\pi(\lambda^{\dagger}),$ (48)
which we estimate separately. When the expectation is not tractable, MCEM algorithms replace $Q$ with an estimator thereof. In this instance, the expectation is taken precisely with respect to $\pi(v_{r},z,y|v_{s},\theta,\lambda)$, for which we developed a sampling algorithm in Section 3.2. Therefore, we can simulate a Markov chain with stationary distribution $\pi(v_{r},z,y|v_{s},\theta,\lambda)$, generate a sequence $(v_{r}^{(l)},z^{(l)},y^{(l)})$ of $\ell$ samples, and replace $Q_{\Theta}$ with the ergodic average
$\log\pi(\theta^{\dagger})+\ell^{-1}\sum_{l=1}^{\ell}\log\pi(V_{R}^{\dagger},Z^{\dagger},v_{s\mathbin{\backslash}\{0\}}|v_{0},y^{\dagger},\theta^{\dagger}).$ (49)
This differs from the Importance sampling approach taken e.g. by gonccalves2023exact, though in this instance, where the data can be highly informative about $Y$, we deem MCMC estimation to be safer. In fact, we require a further estimation step as $\pi(v_{\tau\mathbin{\backslash}\{0\}},z|v_{0},y,\theta)$ is itself intractable. Conveniently, unbiased estimation thereof is easier on the log scale. The log complete transition density is given by
@@
where the path integral can be estimated without bias by uniform subsampling along the path:
@@
¯$\bar{\ell}_{u}(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)=\log h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-(\ddot{\tau}-\dot{\tau})\tilde{\varphi}_{\theta}(z_{u},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),$ (52)
and enrich each Markov chain sample with a sequence $\{u_{(\dot{\tau},\ddot{\tau})}^{(l)}:(\dot{\tau}\sim\ddot{\tau})\in\tau^{(l)}\}$. Assembling those elements, we obtain the unbiased $Q$-estimators
@@ 4.2 M-Step
$\bar{Q}_{\Lambda}(\lambda^{\dagger})=\sum_{i\neq j}\left\lparen\ell^{-1}\sum_{l=1}^{\ell}(n_{ij}^{(l)}\log\lambda_{ij}^{\dagger}+\chi_{i}^{(l)}\lambda_{i}^{\dagger})+(\alpha-1)\log\lambda_{ij}^{\dagger}-\beta\lambda_{ij}^{\dagger}\right\rparen.$ ) . (56)
@@
This solution reveals a limitation of the algorithm: If $\lambda_{ij}^{\dagger}$ is a boundary solution for all $i$, then regime $i$ has 0 probability of being visited under $\pi(y|\lambda^{\dagger})$ and it will be ignored in all subsequent iterations, i.e. the algorithm has absorbing states. We may prevent this behavior by choosing a value $\alpha>1$, but this excludes the pure maximum likelihood case which corresponds to $\alpha=1$, $\beta=0$. If pure ML estimation is required, the absorbing states can be avoided by using a stable algorithm, in the terminology of fort2003convergence. Such an algorithm resets the generator to a safe value when the M-step enters a forbidden, progressively vanishing set.
@@ 4.3 Practical Considerations
As in the posterior sampling case, we deem it helpful to precede the main optimization run with an Euler-approximated run, as sketched out in Supplement C. Since we do not need the algorithm to fully converge at this stage, $\ell$ may be kept constant there. This will typically initiate the main run in fairly close proximity to the MAP, with good tuning parameters for the hidden data update. Tuning parameters are best kept and further adapted across E-steps. In the main run, $\ell$ has to be increased at each E-step to achieve convergence. At the $m$-th E-step, fort2003convergence recommend increasing $\ell_{m}$ at a rate less than exponential, such that $\lim_{m\to\infty}\ell_{m+1}/\ell_{m}=1$. Where large $\ell$ are required to reach convergence, the $Q$-estimate may be constructed from a thinned chain in order to reduce computational burden in the M-step.
@@ 5 Demonstration: Animal Tracking
We proceed with a demonstration of the MCMC and the MCEM algorithms on a moving-resting model for animal movement. The observations were collected from the tagged mountain lion “f109” over multiple years and at irregular intervals, and are included e.g. in the smam package on CRAN. The time series for the lion’s east-west location is shown in Figure 3. Markov switching models have previously been applied e.g. by yan2014moving to capture the alternating moving-resting dynamics in the movement. In yan2014moving, the model takes the form of a tractable 2-regime model, the first of which is a “moving” regime modelled by Brownian motion, and the second is a “resting” regime where the position is fixed. We extend this model by allowing for weak mean reversion, reflecting the fact that the process does not appear to be transient, and we leave all regimes fully symmetrical a priori. Hence, the distinction between “moving” and “resting” states is an entirely implicit property of the posterior, rather than being imposed a priori. The model’s SDE specification is
@@
where the drift function is bounded in absolute value by $\rho_{Y_{t}}\beta_{Y_{t}}$, and time is indexed in hours. For each of the regimes, the diffusion is driven by a separate set of parameters $\theta_{i}=\{\rho_{i},\beta_{i},\mu_{i}\}$. The transition density for this model is intractable, it therefore falls within the scope of our method. We use symmetrical priors for all regimes:
@@ 5.1 MCMC Results
The algorithm converges on a solution with a low volatility “resting” regime, and a high volatility “movement” regime, as presumed by the model of yan2014moving. Figure 1 shows how the resting regime is favored in periods of little observed movement, and vice versa, and how the predictive distribution of $V_{t}$ naturally adjusts. Within those regimes, we observe adequate mixing of $\Theta$, as seen in Figure 4. Since the marginals of $\beta_{0},\beta_{1}$ and $\kappa_{0},\kappa_{1}$ overlap, the respective latent states are most strongly identified by $\rho_{0},\rho_{1}$, resulting in stronger posterior dependence on $Y$ and slower mixing of the volatility parameters.
@@ 5.2 MCEM Results
The algorithm converges on similar “moving” and “resting” regimes as in the posterior sampling case. The estimate of the evidence lower bound (ELBO) stabilizes towards the end of the run, as do the parameter estimates shown in Figure 5. We observe that the parameter estimates coincide with the location of the modes in Figure 6, as estimated from the MCMC output.
@@ 5.3 Approximation Bias
In order to examine the practical impact of approximation biases in Euler-approximated algorithms, we also devise an approximate MCMC algorithm which we outline in Supplement C. This algorithm imputes additional observations at a given rate, with higher imputation rates raising computational cost, but giving a closer approximation to the exact posterior. In the limit of higher imputation rates, the algorithm targets the exact posterior. Indeed, under the hood, the exact algorithm imputes stochastically rather than deterministically, with a rate that tends to increase with Euler approximation bias.
In Figure 6 we compare the posterior marginals of $\theta$ as obtained by the exact algorithm as well as approximate algorithms with different imputation rates. We run the approximate algorithms for 50000 iterations, a target acceptance probabitility of .234 and with imputation rates 0 and 4 per unit time, respectively. The figure indicates that in this instance, doing no imputation results in a sizable bias, while an imputation rate of 4 per unit time is sufficient to eliminate most of that bias. In our implementation, the approximate run with imputation rate 4 runs about 5 times as fast as the exact run. It also has a slightly higher statistical efficiency since it uses the more efficient Metropolis-Hastings, rather than Portkey Barker. We expect that an optimized implementation of the EA3-algorithm would close that performance gap. Moreover, in the absence of the exact benchmark, further imputation would be required to ensure that most of the bias has been eliminated.
@@ 6 Simulation Study
where $\rho$ is a scale parameter, $\beta$ is the reproduction rate and $\kappa^{-1}$ is the carrying capacity of the environment. We also set symmetrical priors for all regimes:
@@
Hence, the same factorization over parameter sets $\theta_{i}=\{\rho_{i},\beta_{i},\kappa_{i}\}$ arises as in Section 5. We follow the efficiency notion of average CPU seconds per effective sample (S/ES), and estimate it from the output of the MCMC algorithm. Both the average seconds per iteration (S/I) and the average number of iterations per effective sample (I/ES) are estimated from MCMC output for various statistics. The computational cost is part deterministic and part random, with either part affected differently in the scaling regimes. The deterministic part of the cost per iteration is linear in the number of observations in both regimes. For the outfill regime, the optimistic scenario is that random costs remain linear in expectation, while the effective sample size remains constant. For the infill regime, we note that random costs depend on the length of the time series and the uncertainty about the diffusion bridges. Since the length of the series is constant but uncertainty is reduced, random costs should decrease. Conclusions from those experiments have limited external validity, and should be seen as setting a benchmark for the behavior of the algorithms under favorable circumstances, i.e. for models that are fairly smooth in $\theta$ and exhibit sufficient posterior concentration rates. We use the integrated autocorrelation time estimator for effective sample size estimation, as seen e.g. in sokal1997monte.
@@
We obtain estimates from MCMC runs of 10000 iterations for each regime, where we’ve discarded 1000 iterations for adaptation, and preceded the main run by 1000 iterations of the approximate algorithm. The chain is thinned to every 10th sample to speed up summary computation. We target an acceptance probability of .2, and apply a “portkey” setting of $\omega^{-1}$ in the parameter update and .01 in the hidden data update. We use one divide-and-conquer split for the “base” and “infill” designs, and two splits for the “outfill” design, which follows the $\mathcal{O}(\log_{4}\omega)$ scaling recommended in stumpf2024bf.
We show performance metrics in Figure 8. The (S/I) measurements show a long tail in iteration time, though we avoid a substantial increase in outliers in the “outfill” regime. The (I/ES) measurements indicate that we meet our objective of achieving stable mixing across the designs. In the outfill regime, the (S/ES) measurements are in line with the $n\log n$ cost scaling observed by stumpf2024bf. In the infill regime (S/ES) is sublinear n $|s|/\omega$ due to tigher bounds on the latent diffusion process.

### v2

@@ 3.4 Generator Update
The update to $\pi(\lambda|y)\propto\pi(y|\lambda)\pi(\lambda)$ does not involve the augmented transition densities, it is therefore the easiest update to carry out. For jump times $r$, the density $\pi(y|\lambda)$ with respect to $\mathbb{L}$ is given by
@@
[…]
@@ 3.5 Practical Considerations
In diffusion settings, the main practical challenge that Bernoulli factory MCMC faces is proposal sensitivity - the notion that the 2-coin algorithm runtime can explode when proposing a move to some regions of the state space, especially those associated with large diffusion drift. This is due to the fact that in such regions, the discrepancy between the diffusion measure and the Brownian bridge proposals is large, and the success probability of Poisson coins is low. Under a correctly specified model, the diffusion path avoids regions of $\mathcal{V}$ with high drift, and conversely, parameters associated with large drift usually have low posterior probability. Accordingly, the $\theta$-proposals in particular need to be carefully chosen. We take two steps to mitigate this issue, the first of which is to apply the Portkey Barker algorithm of [45], which truncates the infinite 2-coin loop at the cost of lesser MCMC efficiency. In practice, the main advantage of this modification is that it discards proposals with potentially very low acceptance probability more quickly. Moreover, the Portkey approach can be extended to the divide-and-conquer Bernoulli factory. The other measure consists of initializing the chain and pre-adapting the proposals by running a chain on an Euler-approximated posterior first. We specify the approximation in Supplement C. This allows us to then start the exact chain at a better location in the state space, with reasonable tuning parameters. We adopt both mitigation strategies in Sections 5 and 6.
More specific challenges apply to Markov switching diffusions, which can exhibit large drift even under the correct specification. This occurs e.g. when switching between two regimes with different stationary means, upon which the process may experience strong drift towards the new equilibrium. This is partially mitigated by refining the bounds on the diffusion path, as recommended by [22], which increases the success probability of Poisson coins. The shortcoming of this intervention is the difficulty of tuning it adaptively during the MCMC run, and usually the chain has to be fully reset to modify the extent of the refinement. We also note that fairly strong posterior dependence between $\Theta$ and $Y$ can arise, which negatively affects mixing of the Gibbs chain. This tendency, while not catastrophic, is observed in the experiment in Section 5. Indeed, the corresponding model is invariant to label permutations, and the Markov chain will typically only visit one of the posterior’s equivalent modes. Moreover, it is common for the chain to drop one of the states during the transient phase, upon which the update for the parameters corresponding to that state may walk randomly. Therefore, during adaptation, we set the parameters of any inactive state equal to the parameters of an active state, which results in quick re-introduction of the inactive state.
@@ 4 Exact MAP and Maximum Likelihood Estimation
[…]
The MAP estimator also corresponds to the maximum likelihood estimator when setting $\pi(\theta,\lambda)\propto 1$. In this section, we adapt an approach originally proposed for Ito diffusions in [8]. It consists of constructing a Monte Carlo EM algorithm, which alternates between an approximate E-step, where we construct a Monte Carlo estimator $\bar{Q}$ of the function
[…]
@@
One benefit of the MCEM approach is that there is a substantial methodological overlap with posterior sampling, and we can make use of the hidden data update of Section 3.2 in devising the E-step. The MCEM algorithm also exploits the same conditional independence structure. Moreover, because it merely requires an unbiased estimate of the log-likelihood, the MCEM algorithm avoids some of the scaling issues with the MCMC parameter update, discussed in Section 3.3. Conversely, even an “exact” MCEM algorithm is not guaranteed to converge to the global maximum, though convergence to a local maximum can be proven in some circumstances [18].
We discuss the construction of $\bar{Q}$ in Section 4.1 before proceeding with its maximization in Section 4.2, and close out with a discussion of various implementation aspects in Section 4.3.
@@ 4.1 E-Step
[…]
@@
$Q_{\Lambda}(\lambda^{\dagger},\lambda)=\operatorname{\mathrm{E}}_{V_{R},Z,Y}\left[\log\pi(Y|\lambda^{\dagger})|v_{s},\theta,\lambda\right]+\log\pi(\lambda^{\dagger}).$ (46)
When the expectation is not tractable, MCEM algorithms replace $Q$ with an estimator thereof. In this instance, the expectation is taken with respect to $\pi(v_{r},z,y|v_{s},\theta,\lambda)$, for which we developed a sampling algorithm in Section 3.2. Therefore, we can simulate a Markov chain with stationary distribution $\pi(v_{r},z,y|v_{s},\theta,\lambda)$, generate a sequence $\{(v_{r}^{(l)},z^{(l)},y^{(l)})\}$ of $\ell$ samples, and replace $Q_{\Theta}$ with the ergodic average
[…]
@@
where the time integral can be estimated without bias by uniform subsampling along the path:
@@
$\bar{\ell}_{u}(z_{(\dot{\tau},\ddot{\tau})},v_{\ddot{\tau}}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)=\log h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})-(\ddot{\tau}-\dot{\tau})\tilde{\varphi}_{\theta}(z_{u},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),$ (50)
and enrich each Markov chain sample with a sequence $\{u_{(\dot{\tau},\ddot{\tau})}^{(l)}:(\dot{\tau}\sim\ddot{\tau})\in\tau^{(l)}\}$. Notice how the variance of this estimator is easily controlled, and how this differs from an MCEM algorithm that further augments the complete likelihood to obtain a tractable $Q$-estimator, which is likely to be slower to converge due to increased dependency between the E- and M-steps.
Assembling those elements, we obtain the unbiased $Q$-estimators
@@ 4.2 M-Step
[…]
@@
This solution reveals a limitation of the algorithm: If $\lambda_{ij}^{\dagger}$ is a boundary solution for all $i$, then regime $i$ has 0 probability of being visited under $\pi(y|\lambda^{\dagger})$ and it will be ignored in all subsequent iterations, i.e. the algorithm has absorbing states. We may prevent this behavior by choosing a value $\alpha>1$, but this excludes the pure maximum likelihood case which corresponds to $\alpha=1$, $\beta=0$. If pure ML estimation is required, the absorbing states can be avoided by using a stable algorithm, in the terminology of [18]. Such an algorithm resets the generator to a safe value when the M-step enters a forbidden, progressively vanishing set.
@@ 4.3 Practical Considerations
As in the posterior sampling case, we deem it helpful to precede the main optimization run with an Euler-approximated run, as sketched out in Supplement C. Since we do not need the algorithm to fully converge at this stage, $\ell$ may be kept constant there. This will typically initiate the main run in fairly close proximity to the MAP, with good tuning parameters for the hidden data update. Tuning parameters are best kept and further adapted across E-steps. In the main run, $\ell$ has to be increased at each E-step to achieve convergence. At the $m$-th E-step, [18] recommend increasing $\ell_{m}$ at a rate less than exponential, such that $\lim_{m\to\infty}\ell_{m+1}/\ell_{m}=1$. Where large $\ell$ are required to reach convergence, the $Q$-estimate may be constructed from a thinned chain in order to reduce computational burden in the M-step.
@@ 5 Demonstration: Animal Tracking
In this section, we demonstrate the MCMC and the MCEM algorithms on a moving-resting model for animal movement. The observations were collected from the tagged mountain lion “f109” over multiple years and at irregular intervals, and are included e.g. in the smam package on CRAN. The time series for the lion’s east-west location is shown in Figure 3. Markov switching models have previously been applied e.g. by [46] to capture the alternating moving-resting dynamics in the movement. In [46], the model takes the form of a tractable 2-regime model, the first of which is a “moving” regime modelled by Brownian motion, and the second is a “resting” regime where the position is fixed. We extend this model by allowing for weak mean reversion, reflecting the fact that the process does not appear to be transient, and we leave all regimes fully symmetrical a priori. Hence, the distinction between “moving” and “resting” states is an entirely implicit property of the posterior, rather than being imposed a priori. The model’s SDE specification is
@@
where the drift function is bounded in absolute value by $\rho_{Y_{t}}\beta_{Y_{t}}$, and time is indexed in hours. For each of the regimes, the diffusion is driven by a separate set of parameters $\theta_{i}=\{\rho_{i},\beta_{i},\mu_{i}\}$. For a given regime, the stationary density is $\pi(a)\propto\operatorname{\mathrm{cosh}}^{-2\beta_{i}/\rho_{i}}\left[a-\mu_{i}\right]$, as obtained by finding the stationary solution of the Fokker-Planck equation, which is similar to a Laplacian density. The transition density for this model is intractable, it therefore falls within the scope of our method. We use symmetrical priors for all regimes:
@@ 5.1 MCMC Results
The algorithm converges on a solution with a low volatility “resting” regime, and a high volatility “movement” regime, as presumed by the model of [46]. Figure 1 shows how the resting regime is favored in periods of little observed movement, and vice versa, and how the predictive distribution of $V_{t}$ naturally adjusts. Within those regimes, we observe adequate mixing of $\Theta$, as seen in Figure 4. Since the marginals of $\beta_{0},\beta_{1}$ and $\kappa_{0},\kappa_{1}$ overlap, the respective latent states are most strongly identified by $\rho_{0},\rho_{1}$, resulting in stronger posterior dependence on $Y$ and slower mixing of the volatility parameters.
@@ 5.2 MCEM Results
The algorithm converges on similar “moving” and “resting” regimes as in the posterior sampling case. The estimate of the evidence lower bound (ELBO) stabilizes towards the end of the run, as do the parameter estimates shown in Figure 5. The parameter estimates coincide with the location of the modes in Figure 6, as estimated from the MCMC output.
@@ 5.3 Approximation Bias
In order to examine the practical impact of approximation biases in Euler-approximated algorithms, we also devised an approximate MCMC algorithm which we outline in Supplement C. This algorithm imputes additional observations at a given rate, with higher imputation rates raising computational cost, but giving a closer approximation to the exact posterior. In the limit of higher imputation rates, the algorithm targets the exact posterior. Indeed, under the hood, the exact algorithm imputes stochastically rather than deterministically, with a rate that tends to increase with Euler approximation bias.
In Figure 6 we compare the posterior marginals of $\theta$ as obtained by the exact algorithm as well as approximate algorithms with different imputation rates. We run the approximate algorithms for 50000 iterations, a target acceptance probability of .234 and with imputation rates 0 and 4 per unit time, respectively. The figure indicates that in this instance, doing no imputation results in a sizable bias, while an imputation rate of 4 per unit time is sufficient to eliminate most of that bias. In our implementation, the approximate run with imputation rate 4 runs about 5 times as fast as the exact run. It also has a slightly higher statistical efficiency since it uses the more efficient Metropolis-Hastings, rather than Portkey Barker. We expect that an optimized implementation of the EA3-algorithm would close that performance gap. Moreover, in the absence of the exact benchmark, further imputation would be required to ensure that most of the bias has been eliminated.
@@ 6 Simulation Study
where $\rho$ is a scale parameter, $\beta$ is the reproduction rate and $\kappa$ is the carrying capacity of the environment. For $2\beta_{i}/\rho_{i}>1$, the corresponding stationary distribution for a fixed regime is $\operatorname{\mathrm{Gamma}}\left[2\beta_{i}/\rho_{i}-1,2\beta_{i}/(\rho_{i}\kappa_{i})\right]$. We also set symmetrical priors for all regimes:
@@
Hence, the same factorization over parameter sets $\theta_{i}=\{\rho_{i},\beta_{i},\kappa_{i}\}$ arises as in Section 5. We follow the efficiency notion of average CPU seconds per effective sample (S/ES), and estimate it from the output of the MCMC algorithm. Both the average seconds per iteration (S/I) and the average number of iterations per effective sample (I/ES) are estimated from MCMC output for various statistics. The computational cost is part deterministic and part random, with either part affected differently in the scaling regimes. The deterministic part of the cost per iteration is linear in the number of observations in both regimes. For the outfill regime, the optimistic scenario is that random costs remain linear in expectation, while the effective sample size remains constant. For the infill regime, we note that random costs depend on the length of the time series and the uncertainty about the diffusion bridges. Since the length of the series is constant but uncertainty is reduced, random costs should decrease. Conclusions from those experiments have limited external validity, and should be seen as setting a benchmark for the behavior of the algorithms under favorable circumstances, i.e. for models that are fairly smooth in $\theta$ and exhibit sufficient posterior concentration rates. We use the integrated autocorrelation time estimator for effective sample size estimation, as seen e.g. in [43].
@@
We obtain estimates from MCMC runs of 10000 iterations for each regime, where we’ve discarded 1000 iterations for adaptation, and preceded the main run by 1000 iterations of the approximate algorithm. The chain is thinned to every 10th sample to speed up summary computation. We target an acceptance probability of .2, and apply a “portkey” setting of $\omega^{-1}$ in the parameter update and .01 in the hidden data update. We use one divide-and-conquer split for the “base” and “infill” designs, and two splits for the “outfill” design, which follows the $\mathcal{O}(\log_{4}\omega)$ scaling recommended in [44].
We show performance metrics in Figure 8. The (S/I) measurements show a long tail in iteration time, though we avoid a substantial increase in outliers in the “outfill” regime. The (I/ES) measurements indicate that we meet our objective of achieving stable mixing across the designs. In the outfill regime, the (S/ES) measurements are in line with the $n\log n$ cost scaling observed by [44]. In the infill regime (S/ES) is sublinear n $|s|/\omega$ due to tigher bounds on the latent diffusion process.